OpenStudy (anonymous):

how can i find a point that lies on the line y-6=3(x-5)

OpenStudy (whpalmer4):

Plug in a value for \(x\), and solve for the value of \(y\). Then \((x,y)\) is a point on the line. For example, if you had the equation \[y+3 = 2(x+4)\]and you felt like finding the point that was on the line at \(x=13\), you plug in \(x=13\) and find the value of \(y\) that results: \[y+3=2(13+4)\]\[y+3=2*17\]\[y+3=34\]\[y+3-3=34-3\]\[y=31\]So \((13,31)\) is a point on my line.

OpenStudy (anonymous):

okay ill try :) thanks for helping

OpenStudy (whpalmer4):

What is it that Yoda says? :-)

OpenStudy (anonymous):

i filled x in with 10 and i did the math and i got (10,21) for a point on the line

OpenStudy (whpalmer4):

Excellent! https://www.youtube.com/watch?v=BQ4yd2W50No

OpenStudy (anonymous):

is there any chance you can help me find a slope of a line?

OpenStudy (whpalmer4):

I suspect I can.

OpenStudy (anonymous):

so basically the same layout its y-4=-9(x-8)

OpenStudy (anonymous):

im not sure but is the slope -9?

OpenStudy (whpalmer4):

Okay. This form is called point-slope form: \[y-y_1 = m(x-x_1)\]where \(m\) is the slope, and \((x_1,y_1)\) is a point that the line goes through.

OpenStudy (whpalmer4):

so yes, you are correct.

OpenStudy (anonymous):

oh thank you :)

OpenStudy (whpalmer4):

You're welcome!

OpenStudy (anonymous):

i also gave you medal :)

OpenStudy (whpalmer4):

Tanks in advance, film at 11...

OpenStudy (whpalmer4):

(of course, it works better if I say that before you do something for me, but...)